OpenStudy (anonymous):

A parallelogram has sides 17.3 m and 43.4 m long. The height corresponding to the 17.3-m base is 8.7 m. Find the height, to the nearest tenth of a meter, corresponding to the 43.4-m base.

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

since the area doesn't change you may use it to find the missing height

OpenStudy (anonymous):

how? sorry I'm really confused

OpenStudy (anonymous):

 you surely know that area=B*H where B is one of the sides and H the height relative to it. so S1*H1=S2*H2 since the area stays constant